# Roles of SN and AGN Feedback in Shaping the Baryonic Content in a Wide Range of Dark Matter Halo Mass

### v2

@@ 3.3 FEGA25 vs Simulations
Figure 7: Baryon fraction normalized to the universal value as a function of halo mass at $z=0$, as predicted by AGNeject1 (left panel) and AGNeject2 (right panel). Model predictions are compared with three hydrodynamical simulations: EAGLE (Schaye et al. 2015, red), IllustrisTNG (Pillepich et al. 2018, green), and SIMBA (Davé et al. 2019, brown). While neither of our models reproduces the simulation results across the entire halo mass range, AGNeject2 exhibits a qualitative similarity with IllustrisTNG and SIMBA: all three show a distinct “cavity” in the baryon fraction, albeit with varying depth and located at different halo masses. A notable feature of the simulations is the rise of the baryon fraction toward the universal value on cluster scales, whereas our models continue to predict a significant baryon deficit (about 25%) even at the high-mass end, indicating that a substantial amount of baryons remains outside the virialized region.
We end our analysis with a full comparison between our models and predictions from hydrodynamic simulations. Ayromlou et al. (2023) convincingly demonstrated that the EAGLE, IllustrisTNG, and SIMBA simulations all predict different baryon fractions as a function of halo mass, emphasizing that the baryonic content of halos is highly sensitive to the specific physical processes implemented in each model.
@@
A similar comparison between observations and simulations was recently presented by Popesso et al. (2024). Using the public Early Data Release (EDR) eROSITA event file of the eFEDS field (Brunner et al., 2022), they compared the hot gas fraction within $R_{500}$ with predictions from a wide array of simulations, including: BAHAMAS (Salcido et al., 2023), FLAMINGO (Schaye et al., 2023), IllustrisTNG, MAGNETICUM, SIMBA, and MillenniumTNG (Pakmor et al., 2023). Their findings indicate that IllustrisTNG and MillenniumTNG—both of which share similar physical models—tend to overpredict the observed hot gas fractions at nearly all halo masses. On the other hand, FLAMINGO and BAHAMAS, which implement similar AGN feedback schemes, are more consistent with observations. Among all the simulations, only MAGNETICUM reproduces the observed relation remarkably well, while SIMBA does so within the $1\sigma$ scatter, albeit with somewhat lower agreement. Notably, MAGNETICUM—the most observationally consistent model—does not exhibit a cavity as seen in some other simulations and in our model AGNeject2, but instead follows a power-law trend, similar to that observed.
To conclude, the comparison of our models with both simulations and observations shows a broad range of behaviors. Some simulations—as well as our model AGNeject2—predict a cavity in the baryon and hot gas fractions, while others—including our model AGNeject1—do not. Current observations tend to support a power-law behavior in both relations. However, as emphasized above, probing scales below $\log M_{\rm{halo}}\sim 12$ is essential before making definitive conclusions about the validity of the physical prescriptions used in simulations and semianalytic models.
@@ 4 Conclusions
In our state-of-the-art semi-analytic model, FEGA25, we have implemented an additional mechanism—alongside SN feedback—for ejecting hot gas beyond the virial radius of halos. This mechanism is directly associated with AGN feedback and is realized in two different forms: the AGNeject1 model, identical to the one introduced in Contini et al. (2025), and the AGNeject2 model, which differs in strength and becomes primarily active at $z<1$.
@@
Our results also highlight the relative importance of SN and AGN feedback across different halo mass scales. In line with recent hydrodynamic simulation results at $z=0$—which we extend to higher redshifts—we find that SN feedback dominates for $\log M_{\rm{halo}}\lesssim 12.5$, while AGN feedback becomes increasingly significant at higher halo masses. On group and cluster scales, both feedback mechanisms contribute comparably. However, at higher redshifts, the models diverge: AGNeject1 maintains a nearly constant division between SN and AGN feedback, whereas AGNeject2 exhibits significant evolution, with SN feedback remaining dominant down to $z=1$, at all scales.
When compared to both simulations and observations, the two models show differing levels of agreement depending on the specific regime. AGNeject1 is more consistent with observational data, as it predicts a smooth, monotonic increase in baryon and hot gas fractions with halo mass. In contrast, AGNeject2 reproduces the cavity seen in several simulations. Nonetheless, we do not discard either model. The lack of observational data at small halo masses—particularly below $\log M_{\rm{halo}}\sim 12.5$—where AGNeject2 and several simulations predict a re-rise in the gas fractions, coupled with the significant observational scatter at all mass scales, makes it premature to rule out any model. In order to assess the physical plausibility of the cavity, and thus constrain the most realistic feedback mechanism, more observational data are needed below $\log M_{\rm{halo}}\sim 12.$–12.5, with significantly reduced scatter (ideally at least a third to half of the current levels).
@@ Acknowledgements
E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (RS-2025-00514475). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (RS-2022-NR070872). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT. This work was also partially supported by the Institut de Physique des deux infinis of Sorbonne Université and by the ANR grant ANR-19-CE31-0017 of the French Agence Nationale de la Recherche.
@@
Codes and galaxy catalogs used in this study can be obtained by contacting the corresponding author.

### v4

@@ 3.3 FEGA25 vs Simulations
Figure 7: Baryon fraction normalized to the universal value as a function of halo mass at $z=0$, as predicted by AGNeject1 (left) and AGNeject2 (right). Predictions are compared with the EAGLE (J. Schaye et al. 2015, red), IllustrisTNG (A. Pillepich et al. 2018, green), and SIMBA (R. Davé et al. 2019, brown) simulations. Neither model matches the simulations across the full halo mass range, but AGNeject2 resembles IllustrisTNG and SIMBA by producing a distinct “cavity” in the baryon fraction, though with different depth and location. Unlike the simulations, which rise toward the universal fraction on cluster scales, our models still predict a $\sim$25% baryon deficit, implying a substantial reservoir of baryons remains beyond the virialized region.
We end our analysis with a full comparison between our models and predictions from hydrodynamic simulations. M. Ayromlou et al. (2023) convincingly demonstrated that the EAGLE, IllustrisTNG, and SIMBA simulations all predict different baryon fractions as a function of halo mass, emphasizing that the baryonic content of halos is highly sensitive to the specific physical processes implemented in each model.
@@
This feature of the rapid development of the cavity in halos with $12\lesssim\log M_{\rm halo}\lesssim 13$ between $z=1$ and the present time has a physical interpretation. In our AGNeject2 model, this late-time behavior arises because the hot gas ejection mode remains inefficient at high redshift, when the supply of cosmological accretion is still strong and AGN feedback energy is largely spent counteracting cooling flows. After $z\sim 1$, however, two conditions change simultaneously: (i) the accretion of fresh gas slows down as halos transition toward quasi-hydrostatic equilibrium, and (ii) BHs have grown sufficiently massive that the cumulative AGN energy input surpasses the binding energy of the hot atmospheres. This means that, in this time and halo mass regimes, the excess energy is no longer required to offset cooling but can instead drive substantial amounts of gas beyond the virial radius.
This transition explains why AGNeject2 is almost inactive for the first $\sim 7$ Gyr of cosmic time, yet becomes suddenly powerful in expelling hot gas at later epochs. The contrast with AGNeject1, which steadily ejects gas throughout cosmic history, highlights that the two implementations probe different physical regimes: a gradual, continuous removal versus a delayed but more explosive release once AGN feedback overtakes the capacity of halos to retain gas. As a result, AGNeject2 naturally produces the fast depletion of gas, a trend that qualitatively matches the cavity hinted by some observational studies and hydrodynamic simulations, as seen in the analysis above.
A full comparison between observations and simulations was recently presented by P. Popesso et al. (2024), who used the public eROSITA Early Data Release event file of the eFEDS field (H. Brunner et al., 2022) to compare the hot gas fraction within $R_{500}$ with several simulations: BAHAMAS (J. Salcido et al., 2023), FLAMINGO (J. Schaye et al., 2023), IllustrisTNG, MAGNETICUM, SIMBA, and MillenniumTNG (R. Pakmor et al., 2023). They found that IllustrisTNG and MillenniumTNG generally overpredict hot gas fractions, while FLAMINGO and BAHAMAS—with similar AGN feedback—are closer to the data. Only MAGNETICUM reproduces the observed relation very well, with SIMBA consistent within $1\sigma$. Interestingly, MAGNETICUM, the best match to the data, does not display a cavity as seen in some simulations and in our model AGNeject2, but instead follows a power-law trend similar to that observed.
Overall, the comparison shows a wide variety of behaviors. Some simulations, as well as our model AGNeject2, predict a cavity in baryon and hot gas fractions, while others—including our model AGNeject1—do not. Current observations appear to favor a power-law behavior, but probing halos below $\log M_{\rm{halo}}\sim 12$ remains crucial before drawing firm conclusions on the physical models used in simulations and SAMs.
@@ 4 Conclusions
In our state-of-the-art semi-analytic model, FEGA25, we have implemented an additional mechanism—alongside SN feedback—for ejecting hot gas beyond the virial radius of halos. This mechanism is directly associated with AGN feedback and is realized in two different forms: the AGNeject1 model, identical to the one introduced in E. Contini et al. (2025), and the AGNeject2 model, which differs in strength and becomes primarily active at $z<1$.
@@
Our results also highlight the relative importance of SN and AGN feedback across different halo mass scales. In line with recent hydrodynamic simulation results at z=0z=0—which we extend to higher redshifts—we find that SN feedback dominates for $\log M_{\rm{halo}}\lesssim 12.5$, while AGN feedback becomes increasingly significant at higher halo masses. On group and cluster scales, both feedback mechanisms contribute comparably. However, at higher redshifts, the models diverge: AGNeject1 maintains a nearly constant division between SN and AGN feedback, whereas AGNeject2 exhibits significant evolution, with SN feedback remaining dominant down to $z=1$, at all scales.
When compared to both simulations and observations, the two models show differing levels of agreement depending on the specific regime. AGNeject1 is more consistent with observational data, as it predicts a smooth, monotonic increase in baryon and hot gas fractions with halo mass. In contrast, AGNeject2 reproduces the cavity seen in several simulations. Nonetheless, we do not discard either model. The lack of observational data at small halo masses—particularly below $\log M_{\rm{halo}}\sim 12.5$—where AGNeject2 and several simulations predict a re-rise in the gas fractions, coupled with the significant observational scatter at all mass scales, makes it premature to rule out any model. In order to assess the physical plausibility of the cavity, and thus constrain the most realistic feedback mechanism, more observational data (besides those quoted above) are needed below $\log M_{\rm{halo}}\sim 12.$–12.5, with significantly reduced scatter (ideally at least a third to half of the current levels).
@@
An interesting extension of this work concerns the cosmological applications of baryon fraction studies. Accurate modeling of the baryon content in halos is crucial for cluster cosmology and has recently been linked to new probes such as fast radio bursts, where the circum galactic medium gas fraction enters the interpretation of dispersion measures (Z. J. Zhang et al., 2025; Y. Liu et al., 2025). Our SAM, with its flexible feedback prescriptions, can therefore provide predictions directly relevant for these emerging cosmological tests.
@@ Acknowledgements
The authors thank the referee for his very constructive comments which helped to improve the manuscript. E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (RS-2025-00514475). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (RS-2022-NR070872). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT. This work was also partially supported by the Institut de Physique des deux infinis of Sorbonne Université and by the ANR grant ANR-19-CE31-0017 of the French Agence Nationale de la Recherche.
@@
The data used in this work, from our semi-analytic model and NewCluster, are available in HARVARD Dataverse: 10.7910/DVN/EGKUF1 (catalog doi: 10.7910/DVN/EGKUF1)